# Regularized Conditional Diffusion Model for Multi-Task Preference Alignment

### v1

@@ 5.1 Setups and Baselines
We conduct experiments on two benchmark datasets, Meta-World Yu et al. (2020) for multi-task scenarios and D4RL Fu et al. (2020) for single-task scenarios. Within evaluations on MetaWorld, we consider two distinct datasets: near-optimal dataset, which comprises the entire experiences obtained during the training of a SAC Haarnoja et al. (2018) agent for each task, and sub-optimal dataset, encompassing only the initial $50\%$ of the replay buffer. More details about datasets and baselines are provided in §B.
Categorically, our baselines encompass three types of approaches: offline preference-based methods, offline reward-based RL methods, and behavior cloning (BC). Within the realm of offline preference-based methods, our selections include: 1) PT Kim et al. (2023), using a transformer network to model reward functions, integrated with RL algorithms; 2) OPRL Shin et al. (2023), which employs ensemble-based reward functions; 3) OPPO Kang et al. (2023), adopting a one-step process to model offline trajectories and preferences, avoiding reward learning. For methods enjoying access to true reward functions, our selection includes: 1) IQL Kostrikov et al. (2022), which performs in-distribution Q-learning and achieves significant performance; 2) MTDiff He et al. (2023), a method leveraging diffusion models in multi-task scenarios. Since many baselines do not consider multiple tasks, we make modifications and mark them with ‘MT--’.
@@ 5.2 Performance Comparison
Environments BC ${}_{\diamondsuit}$ PT OPRL ${}_{\spadesuit}$ ${}_{\clubsuit}$ CAMP (Ours)
@@ 5.3 Visualization
While showing superior performance, does CAMP learn meaningful representations or perform desired conditional generation? In this part, we map trajectory segments to two-dimensional space via T-SNE Van der Maaten and Hinton (2008) visualization and analyze properties of the trajectory encoder and the diffusion model.
@@
Refer to caption (a) Visualization of trajectory representations on MT-10 Refer to caption (b) Visualization of representations on Hopper-medium-expert
@@ 5.4 Analysis on Generalization Ability
To validate the generalization ability, we evaluate generated trajectories guided by $\hat{w}^{*}_{k}\notin\{w_{i}^{*}\}_{i\in[m]}$ that are learned from trajectories of unseen tasks. $\hat{w}^{*}_{k}$ from those new tasks are learned using the same method as described in Section 3.1. We hold the diffusion model trained on MT-10 tasks fixed and assess its performance when faced with unseen representation conditions. We compare its performance with that of MT-BC, MT-IQL, and MTDiff, all of which are trained using the same settings. As depicted in Table 2, our approach exhibits favorable performance on unseen tasks and outperforms baseline methods by a considerable margin. Further analyses are presented in §F.
@@ 6 Conclusion
This paper introduces an aligned and versatile conditional generation approach across multi- and single-task scenarios. Given multi-task preferences, our method can acquire preference representations that differentiate trajectories across tasks and with different returns, with an optimal representation aligning with the best trajectory for each task. Furthermore, we present a mutual information regularized diffusion model conditioned on learned representations, ensuring the alignment between given representations and generated trajectories. Experimental validation demonstrates the favorable performance and generalization ability of our method. Future work may involve using faster sampling methods to enhance algorithm efficiency or extending to fine-tuning foundation models.
@@ A.1 Derivation of ELBO in Equation (1)
Following previous work Luo [2022], Wang et al. [2023a], the above ELBO provides interpretations for each term:
  • The reconstruction term resembles the part of the ELBO of a vanilla variational autoencoder, and can be optimized using Monte Carlo estimates Luo [2022]. In Ho et al. [2020], this term is learned using a separate decoder.
@@ A.2 Derivation of the loss function in Equation (8)
$\mathbb{E}_{q(\tau_{0})}[\log p(\tau_{0})]\geq\mathbb{E}[\mathcal{L}_{\text{elbo}(\tau_{0},w)}]$ $=\mathbb{E}_{q(\tau_{0},\tau_{1})}\big{[}\mathbb{E}_{q_{\psi}(w|\tau_{0})}\big{[}\log p_{\theta}(\tau_{0}|\tau_{1},w)\big{]}\big{]}-\mathbb{E}_{q({\tau_{0}})}\big{[}D_{\rm KL}(q(\tau_{K}|\tau_{0})\|p(\tau_{K}))\big{]}$ (11)
$\qquad-\sum_{k=2}^{K}\mathbb{E}_{q(\tau_{k-1},\tau_{k},\tau_{0})}\big{[}\mathbb{E}_{q_{\psi}(w|\tau_{0})}\big{[}D_{\rm KL}(q(\tau_{k-1}|\tau_{k},\tau_{0})\|p_{\theta}(\tau_{k-1}|\tau_{k},w))\big{]}\big{]}$
$\qquad-\mathbb{E}_{q(\tau_{0})}[D_{\text{KL}}[q_{\psi}(w|\tau_{0})\|p(w)]].$
According to previous work Luo [2022], this optimization problem can be simplified as:
@@
We introduce a lemma to derive the above inequality. Please refer to lemma 5.1 of Chen et al. [2016] for detailed proofs.
@@ B.1 Datasets
Considering this is focused on an offline learning setting, we followed the methodology of MTDiff He et al. [2023], employing a replay buffer during SAC Haarnoja et al. [2018] training as the offline dataset. For each task, we trained an agent using SAC to progress from a random policy toward an expert policy. Additionally, we categorized two distinct datasets: the near-optimal dataset and the sub-optimal dataset, differing in the number of expert trajectories included. The near-optimal dataset comprises all the replay buffer data, totaling 100 million transitions, while the sub-optimal dataset contains only the initial 50% of this data.
Following previous work Kang et al. [2023], Kim et al. [2023], Shin et al. [2023], We collect preference data with a scripted teacher. In particular, for multi-task preferences, we construct intra-task preferences among trajectories by comparing their returns and inter-task preferences among tasks based on the task relevance, as illustrated in Figure 2.
@@
To evaluate our methods on single-task settings, we conduct experiments on the Mujoco locomotion tasks from D4RL benchmarks. Three tasks Halfcheetah, Walker2d, and Hopper are chosen, with three types of datasets, including medium, medium-replay, and medium-expert. We obtain preferences from scripted teachers, following previous work Kim et al. [2023], Kang et al. [2023], Shin et al. [2023].
@@ B.2 Baselines
To make it suitable for multi-task settings, we make two types of modifications to its original version, named ‘MT-OPPO-p’ and ‘MT-OPPO-w’. For ‘MT-OPPO-p’, we add task id inputs for the conditional policy, so that the original policy $\pi(a|s,z)$ can be extended to $\pi(a|s,z,\text{ID})$. For ‘MT-OPPO-w’, we modify OPPO using the same method as our approach, extending its representations to multiple dimensions, so that the policy becomes $\pi(a|s,\textbf{z},\textbf{ID})$ and z is a multi-dimension representation. Following previous work He et al. [2023], we project task IDs to latent variables via a 3-layer MLP.
@@ B.3 Our implementation
  • We employ 200 denoising steps, consistent with previous work Ajay et al. [2023], He et al. [2023].
@@
B.4 Difference from other methods
@@
MTDiff employs diffusion models for modeling large-scale multi-task offline data. Similar to Decision Diffuser, MTDiff adopts classifier-free guidance but introduces prompt learning for both modeling policies and trajectories. Specifically, it utilizes normalized cumulative returns and regards task-relevant information as prompts. By employing these task-specific prompts as conditions, MTDiff distinguishes between different tasks and generates desired trajectories for each specific task. The prompts consist of expert demonstrations in the form of trajectory segments, akin to the approach in PromptDT Hu et al. [2023]. In contrast, our method eliminates the need for expert demonstrations and instead extracts multi-dimensional representations from multi-task preferences. These representations serve as conditions for the diffusion model. Experiments conducted on MetaWorld tasks demonstrate that our method achieves comparable performance to MTDiff, all without the necessity for reward labels or demonstration prompts.
@@
In the realm of offline preference-based reinforcement learning, OPPO models offline trajectories and preferences in a one-step process, circumventing the need for reward modeling. Specifically, OPPO learns an optimal context using a preference modeling objective and subsequently optimizes a contextual policy. During the learning of contexts, OPPO constructs positive and negative samples and uses the triplet loss to optimize the contexts. With these learned contexts, OPPO then develops a conditional policy, akin to DT Chen et al. [2021] and RvS Emmons et al. [2022]. However, our method utilizes the KL loss and the triplet loss to optimize representation distributions, which align with multi-task preferences. Moreover, our approach focuses on conditional diffusion models and the alignment for trajectory generation. It is noteworthy that our method demonstrates more stable performance and excels in multi-task settings, exhibiting successful generalization to unseen tasks. In contrast, OPPO is specifically designed for the single-task setting.
@@ Appendix C Ablation Study
How does the dimension of $w$ affect our method?
When the dimension of $w$ is reduced to 1, the learned representations in our method resemble those of vectorized reward models Vamplew et al. [2022] or distributional rewards Liu et al. [2023]. Conversely, if the dimension is too high, both learning representations and aligning conditional generation become challenging due to the increased complexity of the representation space. We conduct an ablation study with $|w|\in{1,8,16,32,64}$, where $|w|=1$ and $|w|=64$ represent two extremes. As presented in Figure 7, the performance on MT-10 tasks significantly decreases when $|w|=1$, and it exhibits performance degradation when the dimension of representations is too big. This implies that we should adopt a suitable dimension for preference representations, and we choose $|w|=16$ in our experiments.
Is the mutual information regularization critical?
@@ Appendix D Supplementary Results of The Average Performance in MT-10 Tasks
In this section, we present the complete results of our evaluations. All results are obtained over multiple random seeds. All methods are trained with the same data and adopt the same evaluation metric. Tables 4 and 5 elaborate on the success rates on MT-10 tasks given near-optimal and sub-optimal datasets. A high success rate indicates the model’s proficiency in consistently accomplishing the tasks in MT-10 benchmarks.
@@ Appendix E Additional Explanations About The Problem Setting
Appendix F Additional Analysis About The Generalization Ability
F.1 How to obtain $w_{k}^{*}$ for a new task $k$?
@@
F.2 Dataset requirement?
Our learned diffusion model is designed to model the relationship between preference representations and trajectory segments, rather than directly mapping representations to complete trajectories. The preference representations we learn effectively map different trajectory segments into a representation space and then generate trajectory segments based on the conditions of the representation. Therefore, as long as there are some favorable segments within near-optimal trajectories, the learned distribution will involve optimal trajectories, and we can condition the diffusion model on $w^{*}$ to generate optimal trajectory segments. In other words, our approach does not require expert trajectories. In fact, the experimental results in Figure 4 validate that our method performs well on both near-optimal and sub-optimal data. Nevertheless, If the quality of the offline datasets is extremely poor, for example, if they consist entirely of random data, it becomes challenging to derive informative preferences. Consequently, we cannot learn a representation space capable of clearly distinguishing the quality of trajectory segments, thereby limiting the generalization ability of the diffusion model. We note that poor performance and generalization on random data is also commonly encountered in offline RL Kostrikov et al. [2022] and offline preference learning methods Kim et al. [2023], Kang et al. [2023].
F.3 What enables the generalization to unseen tasks?

### v2

@@ 5.1 Setups and Baselines
We conduct experiments on two benchmark datasets, Meta-World [18] for multi-task scenarios and D4RL [17] for single-task scenarios. Within evaluations on MetaWorld, we consider two distinct datasets: near-optimal dataset, which comprises the entire experiences obtained during the training of a SAC [50] agent for each task, and sub-optimal dataset, encompassing only the initial $50\%$ of the replay buffer. More details about datasets and baselines are provided in §B.
Categorically, our baselines encompass three types of approaches: offline preference-based methods, offline reward-based RL methods, and behavior cloning (BC). Within the realm of offline preference-based methods, our selections include: 1) PT [15], using a transformer network to model reward functions, integrated with RL algorithms; 2) OPRL [16], which employs ensemble-based reward functions; 3) OPPO [43], adopting a one-step process to model offline trajectories and preferences, avoiding reward learning. For methods enjoying access to true reward functions, our selection includes: 1) IQL [4], which performs in-distribution Q-learning and achieves significant performance; 2) MTDiff [8], a method leveraging diffusion models in multi-task scenarios. Since many baselines do not consider multiple tasks, we make modifications and mark them with ‘MT--’.
@@ 5.2 Performance Comparison
Environments BC IQL PT OPRL OPPO OPPO CAMP (Ours)
@@ 5.3 Visualization
While showing superior performance, does CAMP learn meaningful representations or perform desired conditional generation? In this part, we map trajectory segments to two-dimensional space via T-SNE [51] visualization and analyze properties of the trajectory encoder and the diffusion model.
@@
(a) Visualization of trajectory representations on MT-10
(b) Visualization of representations on Hopper-medium-expert
@@ 5.4 Analysis on Generalization Ability
To validate the generalization ability, we evaluate generated trajectories guided by $\hat{w}^{*}_{k}\notin\{w_{i}^{*}\}_{i\in[m]}$ that are learned from trajectories of unseen tasks. $\hat{w}^{*}_{k}$ from those new tasks are learned using the same method as described in Section 3.1. We hold the diffusion model trained on MT-10 tasks fixed and assess its performance when faced with unseen representation conditions. We compare its performance with that of MT-BC, MT-IQL, and MTDiff, all of which are trained using the same settings. As depicted in Table 2, our approach exhibits favorable performance on unseen tasks and outperforms baseline methods by a considerable margin. Further analyses are presented in §H.
@@ 6 Conclusion
This paper introduces a regularized conditional diffusion model for alignment with preferences in multi-task scenarios. Based on the versatile multi-task preferences, our method acquires preference representations that differentiate trajectories across tasks and with different returns, as well as an optimal representation aligning with the best trajectory for each task. By regularizing exiting diffusion models in RL with mutual information maximization between conditions and generated trajectories, our method can generate desired trajectories by conditioning on the optimal representation for each task, ensuring alignment with preferences. Experimental validation demonstrates the favorable performance and generalization ability of our method. Future work may involve using faster sampling methods to enhance algorithm efficiency or extending to fine-tuning foundation models.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No.62306242) and the Yangfan Project of the Shanghai Municipal Science and Technology (Grant No.23YF11462200).
@@ A.1 Derivation of ELBO in Equation (1)
Following previous work [19, 21], the above ELBO provides interpretations for each term:
  • The reconstruction term resembles the part of the ELBO of a vanilla variational autoencoder, and can be optimized using Monte Carlo estimates [19]. In [2], this term is learned using a separate decoder.
@@ A.2 Derivation of the loss function in Equation (8)
$\mathbb{E}_{q(\tau_{0})}[\log p(\tau_{0})]$ $\geq\mathbb{E}[\mathcal{L}_{\text{elbo}(\tau_{0},w)}]$ (11)
$=\mathbb{E}_{q(\tau_{0},\tau_{1})}\big{[}\mathbb{E}_{q_{\psi}(w|\tau_{0})}\big{[}\log p_{\theta}(\tau_{0}|\tau_{1},w)\big{]}\big{]}$
$-\mathbb{E}_{q({\tau_{0}})}\big{[}D_{\rm KL}(q(\tau_{K}|\tau_{0})\|p(\tau_{K}))\big{]}$
$-\sum_{k=2}^{K}\mathbb{E}_{q(\tau_{k-1},\tau_{k},\tau_{0})}\big{[}\mathbb{E}_{q_{\psi}(w|\tau_{0})}\big{[}D_{\rm KL}(q(\tau_{k-1}|\tau_{k},\tau_{0})\|p_{\theta}(\tau_{k-1}|\tau_{k},w))\big{]}\big{]}$
$-\mathbb{E}_{q(\tau_{0})}[D_{\text{KL}}[q_{\psi}(w|\tau_{0})\|p(w)]].$
According to previous work [19], this optimization problem can be simplified as:
@@
We introduce a lemma to derive the above inequality. Please refer to lemma 5.1 of [29] for detailed proofs.
@@ B.1 Datasets
Considering this is focused on an offline learning setting, we followed the methodology of MTDiff [8], employing a replay buffer during SAC [50] training as the offline dataset. For each task, we trained an agent using SAC to progress from a random policy toward an expert policy. Additionally, we categorized two distinct datasets: the near-optimal dataset and the sub-optimal dataset, differing in the number of expert trajectories included. The near-optimal dataset comprises all the replay buffer data, totaling 100 million transitions, while the sub-optimal dataset contains only the initial 50% of this data.
Following previous work [43, 15, 16], We collect preference data with a scripted teacher. In particular, for multi-task preferences, we construct intra-task preferences among trajectories by comparing their returns and inter-task preferences among tasks based on the task relevance, as illustrated in Figure 2.
@@
To evaluate our methods on single-task settings, we conduct experiments on the Mujoco locomotion tasks from D4RL benchmarks. Three tasks Halfcheetah, Walker2d, and Hopper are chosen, with three types of datasets, including medium, medium-replay, and medium-expert. We obtain preferences from scripted teachers, following previous work [15, 43, 16].
@@ B.2 Baselines
To make it suitable for multi-task settings, we make two types of modifications to its original version, named ‘MT-OPPO-p’ and ‘MT-OPPO-w’. For ‘MT-OPPO-p’, we add task id inputs for the conditional policy, so that the original policy $\pi(a|s,z)$ can be extended to $\pi(a|s,z,\text{ID})$. For ‘MT-OPPO-w’, we modify OPPO using the same method as our approach, extending its representations to multiple dimensions, so that the policy becomes $\pi(a|s,\textbf{z},\textbf{ID})$ and z is a multi-dimension representation. Following previous work [8], we project task IDs to latent variables via a 3-layer MLP.
@@ B.3 Our implementation
  • We employ 200 denoising steps, consistent with previous work [5, 8].
@@
B.4 Difference with other methods
@@
MTDiff employs diffusion models for modeling large-scale multi-task offline data. Similar to Decision Diffuser, MTDiff adopts classifier-free guidance but introduces prompt learning for both modeling policies and trajectories. Specifically, it utilizes normalized cumulative returns and regards task-relevant information as prompts. By employing these task-specific prompts as conditions, MTDiff distinguishes between different tasks and generates desired trajectories for each specific task. The prompts consist of expert demonstrations in the form of trajectory segments, akin to the approach in PromptDT [52]. In contrast, our method eliminates the need for expert demonstrations and instead extracts multi-dimensional representations from multi-task preferences. These representations serve as conditions for the diffusion model. Experiments conducted on MetaWorld tasks demonstrate that our method achieves comparable performance to MTDiff, all without the necessity for reward labels or demonstration prompts.
@@
In the realm of offline preference-based reinforcement learning, OPPO models offline trajectories and preferences in a one-step process, circumventing the need for reward modeling. Specifically, OPPO learns an optimal context using a preference modeling objective and subsequently optimizes a contextual policy. During the learning of contexts, OPPO constructs positive and negative samples and uses the triplet loss to optimize the contexts. With these learned contexts, OPPO then develops a conditional policy, akin to DT [53] and RvS [54]. However, our method utilizes the KL loss and the triplet loss to optimize representation distributions, which align with multi-task preferences. Moreover, our approach focuses on conditional diffusion models and the alignment for trajectory generation. It is noteworthy that our method demonstrates more stable performance and excels in multi-task settings, exhibiting successful generalization to unseen tasks. In contrast, OPPO is specifically designed for the single-task setting.
@@ Appendix C Ablation Study
C.1 How does the dimension of $w$ affect our method?
When the dimension of $w$ is reduced to 1, the learned representations in our method resemble those of vectorized reward models [55] or distributional rewards [56]. Conversely, if the dimension is too high, both learning representations and aligning conditional generation become challenging due to the increased complexity of the representation space. We conduct an ablation study with $|w|\in{1,8,16,32,64}$, where $|w|=1$ and $|w|=64$ represent two extremes. As presented in Figure 7, the performance on MT-10 tasks significantly decreases when $|w|=1$, and it exhibits performance degradation when the dimension of representations is too big. This implies that we should adopt a suitable dimension for preference representations, and we choose $|w|=16$ in our experiments.
C.2 Is the mutual information regularization critical?
@@
C.3 Ablation study on the influence of the number of tasks
We have also conducted an ablation study on the number of tasks $K$ and compared the performance on five tasks. The results including average return and average success rate across tasks are shown below.
Table 4: Ablations on the number of tasks.
Tasks $K$=3 $K$=5 $K$=10
Reach-v2 172.9 1209.4 3625.7
Pick-Place-v2 2.8 2.7 564.8
Window-Open-v2 130.2 1195.6 3188.7
Push-v2 - 32.2 1083.3
Drawer-Open-v2 - 586.2 2304.0
Average Success Rate Across 5 Tasks 0.0% 4.0% 51.2%
From these results, it can be observed that as the number of tasks $K$ increases from 3 to 5 to 10, both the average return and success rate in experiments gradually increase. This suggests that when $K$ is not large, increasing the number of tasks $K$ can enhance the performance of a specific task. We attribute this to the speculation that as the number of tasks increases, the representation learning for a specific task improves. As mentioned in Section 3.1, our method learns a task-specific preference representation and an optimal representation by constructing positive and negative samples, where the negative samples include all trajectories from other tasks. When $K$ increases, the number of negative samples significantly increases, which helps our method learn the preference representation under the given task. Therefore, learning from more tasks leads to better representations of a specific task, resulting in improved performance and accuracy for that particular task.
It is also important to note that when $K$ becomes too large, the learning process for the diffusion model becomes more challenging. This is because the diffusion model needs to fit the trajectory distributions of multiple tasks. Especially when the multi-task trajectory data are low-quality, the difficulty of fitting for the diffusion model increases.
@@ Appendix D Supplementary Results of The Average Performance in MT-10 Tasks
In this section, we present the complete results of our evaluations. All results are obtained over multiple random seeds. All methods are trained with the same data and adopt the same evaluation metric. Tables 5 and 6 elaborate on the success rates on MT-10 tasks given near-optimal and sub-optimal datasets. A high success rate indicates the model’s proficiency in consistently accomplishing the tasks in MT-10 benchmarks.
@@ Appendix E Additional Explanations About The Problem Setting
Appendix F Acceleration Sampling for Diffusion models
While multi-step denoising process in the diffusion model generation is time-consuming, our contribution is orthogonal to those sampling acceleration methods for diffusion models, such as DDIM [57], DPM-solver [58], EDP [59], and can be easily combined with them. In fact, we have incorporated the implementation of the DPM-solver into our method.. This improvement has boosted the inference speed by a factor of 8.4 compared to the previous method. Specifically, on an Nvidia RTX3090, the denoising time per step has been reduced from 15.2ms to 1.8ms. However, our initial results indicate a compromise in the quality of the generated trajectories, likely due to the complexity of generating continuous trajectories compared to image-generation tasks. We will explore further enhancements to balance performance and quality in our future work.
Appendix G Further Analysis About MI Regularization
The conditional diffusion model is not enough to provide alignment
The remarkable success of Stable Diffusion and Midjourney has highlighted the potential of diffusion models to generate desired images given textual conditions. However, when diffusion models are applied to offline reinforcement learning and trajectory modeling, the situation appears different. Previous work on return-conditioned diffusion models [5, 14] generates trajectories by conditioning on target returns, but their conditions are often hyperparameters adjusted for each environment. As shown in Figure 1, our tests with different conditions reveal that the relationship between the generation and the conditions is not as expected. Given a higher target return, the diffusion model cannot generate better trajectories. In fact, this phenomenon has been noted in other related works as well [60]. Therefore, the current conditional diffusion models are insufficient to provide the necessary alignment between preference and trajectory in our setting. This is one of the main problems we aim to solve.
Why using MI regularization?
Inspired by the work in the field of image generation, such as InfoVAE [28], InfoGAN [29], and InfoDiffusion [21], we propose the adoption of mutual information regularization. During the learning phase of the diffusion model, it effectively estimates prior noise, with its posterior estimation conditioned on preference representation. In our debugging phase, we observed that noise estimates from previous works often disregard conditioning, leading to the diffusion model’s inability to effectively learn the conditional distribution $p(\tau|c)$. Therefore, in our work, we imposed a constraint of maximizing mutual information on the diffusion model, ensuring a tighter relationship between the posterior noise estimation and the condition. This guarantees that the noise estimation network can capture information about the condition. Consequently, during the denoising generation phase, the trajectories generated by the diffusion model can well correspond to the condition. Moreover, in response to the inefficiency of the multi-step denoising process in diffusion models, we have introduced a reasonable approximation that simplifies the implementation of regularization. The experimental results also demonstrate that enhancing this connection leads to better outcomes than not doing so.
Appendix H Additional Analysis About The Generalization Ability
H.1 How to obtain $w_{k}^{*}$ for a new task $k$?
@@
H.2 Dataset requirement?
Our learned diffusion model is designed to model the relationship between preference representations and trajectory segments, rather than directly mapping representations to complete trajectories. The preference representations we learn effectively map different trajectory segments into a representation space and then generate trajectory segments based on the conditions of the representation. Therefore, as long as there are some favorable segments within near-optimal trajectories, the learned distribution will involve optimal trajectories, and we can condition the diffusion model on $w^{*}$ to generate optimal trajectory segments. In other words, our approach does not require expert trajectories. In fact, the experimental results in Figure 4 validate that our method performs well on both near-optimal and sub-optimal data. Nevertheless, If the quality of the offline datasets is extremely poor, for example, if they consist entirely of random data, it becomes challenging to derive informative preferences. Consequently, we cannot learn a representation space capable of clearly distinguishing the quality of trajectory segments, thereby limiting the generalization ability of the diffusion model. We note that poor performance and generalization on random data is also commonly encountered in offline RL [4] and offline preference learning methods [15, 43].
H.3 What enables the generalization to unseen tasks?